OpenStudy (anonymous):

Explain how the formula for the area of a regular polygon is derived from that of a triangle. I'm so confused with this.

OpenStudy (ghazi):

if you'll keep adding triangle at different sides of it you'll get regular ..like this polygon

OpenStudy (ghazi):

now summing up the areas of triangle will give you area of regular polygon

hero (hero):

Multiply the numerator of \(\frac{bh}{2}\) by n to get \(\frac{nbh}{2}\) which equals \(\frac{Pa}{2}\) n = number of triangles inside the polygon nb = Perimeter h = a = apothem

hero (hero):

 I know the drawing is poor but what I just attempted to draw is a regular polygon split into triangles